OpenStudy (anonymous):

What is the solution of | - 1 + X | = 5

OpenStudy (anonymous):

\[| - 1 + X | = 5 \rightarrow - 1 + X = \pm5\] i.e. - 1 + X = 5 or - 1 + X = -5 i.e. x= 5+1 or x= 1-5 i.e. x= 6 or x= -4 i.e. x= -4, 6 is the solution of the given problem

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

You are most welcome.